Compute the values of the sums below.
(a)5∑(k=1) (k+ 1)
(b)4∑(j=0) (-2) ^j
(c) 10∑(i=1) 3
(d) 8∑(j=0) ( 2^(j+1) -2^j )
1
Expert's answer
2020-10-15T17:36:27-0400

(a) "5\sum_{k=1}^{+\infty}(k+1)=5\, lim_{n\rightarrow+\infty}\sum_{k=1}^{n}(k+1)=5\, lim_{n\rightarrow+\infty}\frac{(n+2)(n+1)}{2}=+\infty"

(b) "4\sum_{j=0}^{+\infty}(-2)^j=4\, lim_{n\rightarrow+\infty}\sum_{j=0}^{n}(-2)^j\, =lim_{n\rightarrow+\infty}\frac{(-2)^{n+1}-1}{-3}=\infty"

(c) "10\sum_{i=1}^{+\infty}3=+\infty"

(d) "8\sum_{j=0}^{+\infty}(2^{j+1}-2^{j})=8\,\,lim_{n\rightarrow+\infty}\sum_{j=0}^{n}(2^{j+1}-2^{j})="

"=8\,\,lim_{n\rightarrow+\infty}(2^{n+1}-1)=+\infty"